The length of a rectangle is 5 more than twice its width, and the area of the rectangle is 88yd^2. Find the dimensions of the rectangle.

Answer by cleomenius(959)

W = x
L length = 2x + 5
x(2x + 5) = 88
2x^2 + 5x - 88 = 0
(2x - 11)(x + 8)
x = 5.5 yd, the width.
2(5.5) + 5 = 16 yd, the length.
5.5 * 16 = 88 yd^2.
This does check, Cleomenius.